# Constraints to Lorentz violation and ultrahigh-energy electrons in D-foamy space-times

### v1

@@ 2.3 Comparison with analogous models of QED
[…]
@@ 3.1 String/Brane-world inspired D(efect)-foam and electrons
Below, we hope to present the plausible mechanisms that could permit stable propagation of electrons for this type of QG-foam of strings. Actually, for any theory exhibiting subluminal refraction effects of photons in a nontrivial vacuum, we have emphasized in the introduction that, one can avoid the occurrence of the process, $e\rightarrow e\gamma$ in vacuo, if the theory entails the lack of i) a description of the photon propagation via the dispersion relation, or ii) an exact energy–momentum conservation. String models we shall examined below are characterized by these two unusual (and useful) properties, respectively, related with their lack of a local-effective Lagrangian formalism. We must stress at this point, that an oversimplified version of the suggestion that electrons do not Cherenkov radiate in the foam, was first put forward in Ellis:2003ua , but to our knowledge, no detailed description has been presented, especially for the models developed subsequently, i.e., in Ellis:2008gg ; Ellis:2009vq and Li:2023wlo , two specific situations that we are concerned with, thereby calling for some clarifications.
@@
$\frac{M_{s}}{g_{s}n_{D}(z)}\geq 3\times 10^{17}~{}\textrm{GeV},\quad n_{D}(z<10)\simeq\textrm{const.},$. , (26)
which plays the role of the QG mass, [^11] Note that from a modern perspective, where large extra dimensions may be introduced, $M_{s}$ is considered arbitrary. Despite the discrepancies in the bounds placed on $E_{\textrm{LV}\gamma}$ like that provided by GRB 090510 Fermi-LAT:2009grb ; Xiao:2009xe where ad hoc assumptions on the intrinsic time lags at the source have been made, it could indeed be probed with the findings from these analyses Shao:2009bv . The negative source timing (i.e., high energy photons are emitted before softer ones) found solely via data fitting agrees with the cooling nature of GRBs. A recent survey Liu:2024qbt of the brightest ever GRB 221009A data finds direct signals for this preburst stage of GRBs. associated with frequency dispersion for the propagation speed of light, with $n_{D}$ being the D-particle number density, i.e., the number of defects encountered by the photonic matter per string length $\ell_{s}(\equiv M_{s}^{-1})$. Such an explanation appears remarkably consistent also with the TeV spectra of active galactic nuclei (AGNs) Li:2020uef which constitute some other examples of observed effects on retarded photon arrivals, with higher energy photons arriving later. It is also important to note that, the delay effect (25), being independent of photon helicity, is nonbirefringent, evading the otherwise tight bounds from polarized measurements He:2022gyk  (see also ref. Li:2021gah in this regard).
@@
This is consistent with a local effective action proposal for the models Mavromatos:2010ar where a charged fermion $\psi$ has no tree-level coupling to the foam recoil-velocity field $u_{i}$:
@@ 3.2 Absence of any Cherenkov effect
[…]
@@
Next we come to the analysis on the (hard) Cherenkov effect of electrons in the presence of foam. At this stage, we reiterate that the formalism of Li:2023wlo goes beyond the local effective Lagrangians that exact energy–momentum conservation in the process is, by its very nature, based upon, due to the presence of Finsler dynamical background (34), where foam defects act as an external quantum-gravity ‘environment’, resulting in energy violations for particle reactions Ellis:2000sf . As we discussed in Li:2023wlo , for a $1\rightarrow 2$-body decay process (relevant for us here), the induced energy nonconservation in the observable particle subsystem, is associated with the kinetic energy of the heavy D-particle in the recoiling foam: [^13] For a ${\cal N}$-particle interaction, a theorem holds, stating, that the total energy of all the observable particles is in general violated by the requirement of energy conservation for the complete matter $+$ recoiling D-brane system: ±$\prescript{({\cal N})}{}{\sum\nolimits_{\ell}^{\pm}}E_{\ell}\simeq\frac{M_{s}}{2g_{s}}u_{\|({\cal N})}^{2},\quad\prescript{({\cal N})}{}{\sum\nolimits_{\ell}^{\pm}}{\bf p}_{\ell}=M_{D}{\bf u}_{\|({\cal N})},\quad\Rightarrow\quad\delta\overline{E}_{D}^{({\cal N})}\coloneqq\overline{\prescript{({\cal N})}{}{\sum\nolimits_{\ell}^{\pm}}E_{\ell}},$ here $u_{\|({\cal N})}$ is a proper generalization of (34), c.f. ${\bf u}_{\|}=\sum_{\ell}^{\pm}({\bf p}_{\ell}/M_{D})$ and the $\sum^{\pm}$ represents, e.g., in (42): ±$\sum^{\pm}\!E_{\ell}=E_{\textrm{in}}-\sum E_{\textrm{out}}=E_{1}-E_{2}-E_{3}$ as ${\cal N}=3$. For details see appendix A of Li:2023wlo and/or ref. Ellis:2000sf .
@@
$-2m_{e}^{2}=E_{th}\Bigl{(}4\delta\overline{E}_{D}^{(3)}-\frac{g_{s}}{2M_{s}}\mathfrak{d}_{\gamma}^{2}E_{th}^{2}\Bigr{)},\quad E_{th}\coloneqq E_{e,th}^{\textrm{D-foam}}$ (45)
@@
where according to Li:2023wlo we assume this equality to hold. The condition, that the relativistic electron/positron of momentum ${\bf p}$, responsible for producing UHE IC radiation of the Crab Nebula, are near or below the corresponding threshold implied by (45),
@@
$-\frac{g_{s}}{2M_{s}}\bigl{(}2\varsigma_{{\cal I}}{}^{\gamma}-\mathfrak{d}_{\gamma}^{2}\bigr{)}\lvert{\bf p}\rvert^{3}\leq 2m_{e}^{2}.$ (48)
@@
We have to stress that these values imply no extra suppression of the effect, but natural values for both parameters in such brane models where both $\mathfrak{d}_{\gamma}^{2}$ and $\varsigma_{{\cal I}}{}^{\gamma}$ being dimensionless though, are free to vary, independently. In case lower (TeV) $M_{D}$ are considered (as in some low-scale string schemes Antoniadis:1999rm ) the scenario envisaged by the data requires a smaller variance; this can also be naturally realized, with a much more dilute population of defects, provided $\mathfrak{d}_{\gamma}^{2}$ is proportional to the D-particle density, $n_{D}$, which is another free parameter of the bulk string cosmology model of Ellis:2004ay we used here. Although, from the observational requirement one concludes a limit (49) for $\mathfrak{d}_{\gamma}^{2}>2\varsigma_{{\cal I}}{}^{\gamma}$, this is not fine-tuning in our opinion, but implies the ‘recovery’ of special-relativistic kinematics, which is indeed respected by strings, due to underlying Lorentz symmetry that is broken, or better, reduced by the (stringy) QG ground state. Indeed we have noted that the model may be characterized Li:2023wlo via a reduced-Lorentz symmetry (akin to the one found from coupling Ellis:1999sf a fermionic system to the gravitational field (24)), with which the challenges brought about by superluminal neutrino instabilities, are properly resolved. These have not been discussed in our previous study of this isotropic stochastic D-foam, presented in Li:2023wlo , where attention was restricted to neutrino speed with favorable signal of the effect in astrophysical data.
@@ 4 Evasion of Cherenkov and other strict limits on photons
We know that modified dispersions of the type (1) if used to encapsulate QG phenomenology have a broad scope to limit LV. Besides vacuum Cherenkov effect focused on here, other effects regarding photon sector include time-of-flight delay in $\gamma$-ray detections. In Shao:2009bv ; Li:2020uef in particular, it was suggested that, even with the incorporation of intrinsic timing of $\gamma$-rays at the respective source a tiny subluminal speed variation linearly scaling as photon energy could still be a significant contribution to the observed time delay, with an associated scale $E_{\textrm{LV}\gamma}\simeq 3\times 10^{17}$ GeV for the effect $\Delta v_{\gamma}=-E_{\gamma}/E_{\textrm{LV}\gamma}$. If one attempts to interpret it with the corresponding anomalous dispersion relations of the photons, the hard Cherenkov effect is allowed by kinematics as a result. It is in that context that UHE electrons deduced from the LHAASO measurements of the Crab IC spectrum have been claimed to raise the scale well above the Planck energy He:2022jdl ; specifically, the presence of 2.3 PeV long living electrons require the scale to be greater than $1.2\times 10^{25}$ GeV, which exceeds the sensitivity of GRB photon time delay studies to such subluminal photon LV properties Shao:2009bv ; Li:2020uef by about seven orders of magnitude.
@@
    or equivalently, for UHE electrons, $m_{e}\ll p\simeq E\geq E_{th}\equiv(4m_{e}^{2}{\cal E}_{\ast})^{1/3}$ from eq. (47). The new energy scale ${\cal E}_{\ast}$ we introduce here for $e\rightarrow e\gamma$ is given by
@@
    where $\Delta v_{\gamma}=-\mathfrak{d}_{\gamma}^{2}E_{th}/M_{D}=-E_{th}/(2E_{\textrm{LV}\gamma})$ at threshold is used. Thus if the photon-velocity ‘anomaly’ can finally be attributed to this type of foam, indicated by the last equality, the threshold limits as derived He:2022jdl for the case of ($\eta=0$, $\xi<0$), are actually imposed for $\varDelta$ or on the scale ${\cal E}_{\ast}$ in our case, though can be evaded if ${\cal E}_{\ast}\leq 0$ or easily satisfied in cases ${\cal E}_{\ast}>0$ via an assignment for the value of $\varsigma_{{\cal I}}{}^{\gamma}$ as above; they do not constrain the actual subluminal velocity $\Delta v_{\gamma}$ and the corresponding scale $E_{\textrm{LV}\gamma}$. The latter, related to $M_{s}/(\mathfrak{d}_{\gamma}^{2}g_{s})$, can only be limited in relation to time lags in the arrival of cosmic $\gamma$-rays as in eq. (26).
@@
We must emphasize once more, however, that any speculation about Lorentz violations in cosmic radiation needs support from more data. With many more UHE astrophysical $\gamma$-ray data being accrued by LHAASO and other experiments, it might be able to disentangle source from possible propagation effects due to fundamental (string) physics. On the other hand, from the perspective of the D-particle models endorsed here, other cosmic probes like neutrinos may help to discriminate between the above two approaches: the string-stretched linear-in-energy effects (30), (31), when applied to neutrinos, stay the same Ellis:2008gg , while for the models of Li:2023wlo they become entangled with the particle/antiparticle nature of neutrinos as evidenced from eq. (35). A more detailed recapitulation of their respective predictions will be given elsewhere.
@@ 5 Discussion
In this article we have revisited electron Cherenkov effect in vacuo, a nonstandard Lorentz-breaking effect for which we have no evidence till today, and thus, whose absence in particular has been used by some recent analyses He:2022jdl ; Li:2022ugz to improve prior limits on certain types of Lorentz violations in QED. Clearly, the observational hint against such a process is indeed relevant for a pure-kinematics phenomenological formalism (where just modified dispersion is postulated) as adopted therein, but in ways that it seems difficult to quantify, due to the dependence on the strength of interaction (dynamics/matrix element), which is not known in this context. So, here, we have contemplated the implications of these findings from the perspective of the comments offered in the introduction section 1 concerning the plausibility (and, practically, lack thereof) of conspiracies between modifications to kinematics and that of dynamics. We see that the latter is not particularly important as deriving the rates, and it is verified that in cases of our interest, above the threshold the rate is so tremendous that, it would allow a threshold analysis alone with trans-Planckian sensitivity which relies on photon observations from distant nebulae such as the Crab.
@@
However as it has become clear in section 4, one cannot use this result to set constraints to certain models of (stringy) QG, according to which effective subluminal refractive indices of a 4D nonstandard vacuum, as ‘seen’ by the photons, may be induced, via specific types of reduced-Lorentz invariance. In particular, in section 3, we discuss such instances arising in brane-world scenario and we see that within D-foam models hereby provided albeit linearly modified photon propagation is found to characterize the situation electrons and, in general charged excitations, remain undisturbed by the presence of space-time D-brane defects. At first sight this is just the LV parameter configuration of $(\eta=0$, $\xi<0)$, allowed by a general setting (1), pointing likely towards the possibility of hard Cherenkov effect. While in such noneffective field theories for QG foam, this is not the case. In subsection 3.2, we described briefly the reasoning based on two specific foam situations, including the cases of stretched strings and isotropic, stochastic recoil fluctuations. In both cases, although fermions could travel faster than photons with comparable energies, at least effectively no such a radiation happens. Therefore in the contexts of such models one can develop a consistent description for the intriguing result of refs. Shao:2009bv ; Li:2020uef where a linearly Planck scale suppressed light-speed variation was suggested from analyzing travel times of GRB and AGN photons, and which may thus serve as a support to such string models Li:2021gah .
@@
To make our article most useful for the reader who looks for the severest bound ever to this phenomenon, and hence on these two types of modifications in particle dispersions, we note that these scales can be further constrained with LHAASO’s latest report on an UHE $\gamma$-ray bubble LHAASO:2023uhj in the direction of the active star-forming region Cygnus-X, the region from which the previous $\sim 1.4$ PeV photon event was found. Gamma-rays obtained towards the Cygnus bubble are observed up to an unprecedented high energy of 2.48 PeV. [^16] We note that the detected UHE photons spread over the entire bubble region make unlikely the leptonic IC origin of the radiation and indicate the involvement of hadronic process LHAASO:2023uhj , so that the argument leading to electron Cherenkov bound as applied to the Crab is not valid in this case. So, we find the constraints:
@@
The estimates we make tighten the strongest one (among the existing values, say of refs. LHAASO:2021opi ; Li:2021tcw ; He:2022jdl ) by a factor of $\sim(12/7)^{3}\approx 5$. The inclusion of more general cases having $\xi\neq 0$ and $\eta\neq 0$, is straightforward, and, parallels the analyses carried out in He:2022jdl using 1.4 PeV as the greatest photon energy, prior to the discovery of the $\sim 2.5$ PeV event. While an analysis in this line does not present any technical difficulty, it requires a separate investigation which takes us out of the purpose of the present article.
Finally, one should observe that the trans-Planckian lower limits for these types of LV, or equivalently, the small values of the corresponding dimensionless parameters $\xi$ and $\eta$ are consistent with the subluminal feature of the D-foam-induced vacuum refractive indices for photons, capable of accommodating the hint of light-speed variation exposed in the time-of-flight tests Shao:2009bv ; Li:2020uef . Indeed in the context of such brane models, as advocated here, photons are stable (i.e., do not decay), so that Li:2021gah the strong $\gamma$-decay constraints thereof LHAASO:2021opi ; Li:2021tcw ; He:2022jdl , as well as the most stringent result just derived, are inapplicable. For reconciliation of such models with LHAASO data, especially in this respect, we refer the reader to our article Li:2021tcw for details. Moreover, it should be noted that the peculiarity of the transparency of D-foam to electrons or charged probes, for specifically stringy properties outlined above, [^17] It is for the same reason that such models also survive from the severe constraints (e.g., He:2024ljr ) that would be obtained from models with LV modifications in protons and/or pions and thus modifications in hadronic GZK cutoff, which is unaffected due to vanishing or suppressed interactions between particles involved (i.e., protons, or pions whose constituents are charged) and D-foam. provides additional reasons for escaping from the otherwise very tight bounds for putative LV effects on electrons/positrons, including (55). For this latter point, see also Li:2022ugz . This is similar to the way, discussed in that paper, out of electron (soft) Cherenkov limits on this parameter imposed from the positive side.

### v2

@@ 2.3 Comparison with analogous models of QED
$P(E)\equiv-\frac{\textrm{d}E}{\textrm{d}t}\approx\frac{e^{2}}{4\pi}E^{2}\biggl[\Bigl(\frac{7}{24}\kappa-\frac{1}{16}\kappa^{2}+{\cal O}(\kappa^{3})\Bigr)+m_{e}\textrm{-dependent terms}\biggr],$ (22)
@@
We make an important remark here. [^9] We are grateful to an anonymous referee for bringing this issue to our attention and encouraging us to clarify this point for the reader’s benefit. The phase space estimates, which form the basis of our analysis, may not be entirely accurate for determining the rates of Lorentz-violating processes when modifications of particle dispersion relations depend on helicity and particle versus antiparticle identity. This is particularly true for dimension-5 LV QED Myers:2003fd : as shown in Jacobson:2005bg , properly accounting for this dependence complicates matters, generally altering the behavior of the decay rates. For a detailed discussion of the role that spin structure plays in this theory, we refer the reader to ref. Altschul:2010nf . However, as previously emphasized, our explicit calculations here are restricted to spin-independent LV cases. While such modifications are not allowed for photons in local field theory, they do align with the D-brane/string models presented later on.
@@ 3.1 String/Brane-world inspired D(efect)-foam and electrons
Below, we hope to present the plausible mechanisms that could permit stable propagation of electrons for this type of QG-foam of strings. Actually, for any theory exhibiting subluminal refraction effects of photons in a nontrivial vacuum, we have emphasized in the introduction that, one can avoid the occurrence of the process, $e\rightarrow e\gamma$ in vacuo, if the theory entails the lack of i) a description of the photon propagation via the dispersion relation, or ii) an exact energy–momentum conservation. String models we shall consider below are characterized by these two unusual (and useful) properties, respectively, related with their lack of a local-effective Lagrangian formalism. We must stress at this point, that an oversimplified version of the suggestion that electrons do not Cherenkov radiate in the foam, was first put forward in Ellis:2003ua , but to our knowledge, no detailed description has been presented, especially for the models developed subsequently, i.e., in Ellis:2008gg ; Ellis:2009vq and Li:2023wlo , two specific situations that we are concerned with, thereby calling for some clarifications.
@@
$\frac{M_{s}}{g_{s}n_{D}(z)}\geq 3\times 10^{17}~\textrm{GeV},\quad n_{D}(z<10)\simeq\textrm{const.},$ (26)
which plays the role of the QG scale [^12] Note that from a modern perspective, where large extra dimensions may be introduced, $M_{s}$ is considered arbitrary. Despite the discrepancies in the bounds placed on $E_{\textrm{LV}\gamma}$ like that provided by GRB 090510 Fermi-LAT:2009grb ; Xiao:2009xe where ad hoc assumptions on the intrinsic time lags at the source have been made, it could indeed be probed with the findings from these analyses Shao:2009bv . The negative source timing (i.e., high energy photons are emitted before softer ones) found solely via data fitting agrees with the cooling nature of GRBs. A recent survey Liu:2024qbt of the brightest ever GRB 221009A data finds direct signals for this preburst stage of GRBs. associated with frequency dispersion for the propagation speed of light, with $n_{D}$ being the D-particle number density, i.e., the number of defects encountered by the photonic matter per string length $\ell_{s}(\equiv M_{s}^{-1})$. Such an explanation appears remarkably consistent also with the TeV spectra of active galactic nuclei (AGNs) Li:2020uef which constitute some other examples of observed effects on retarded photon arrivals, with higher energy photons arriving later. It is also important to note that, the delay effect (25), being independent of photon helicity, is nonbirefringent, evading the otherwise tight bounds from polarized measurements Li:2025stf ; He:2022gyk  (see also ref. Li:2021gah in this regard).
@@
This is consistent with a local effective action proposal for the model Mavromatos:2010ar in which charged fermion $\psi$ has no tree-level coupling to the foam recoil-velocity field $u_{i}$:
@@ 3.2 Absence of any Cherenkov effect
which represents the effect averaged out by a brane observer. The suffix $\|$, appearing above, indicates components along brane longitudinal dimensions $i=1,2,3$ for the (dimensionless) momentum fraction variable $\zeta_{i\|}=\Delta k_{i}/k_{i}$, which is taken to be Gaussian.
@@
Next we come to the analysis on the (hard) Cherenkov effect of electrons in the presence of foam. At this stage, we reiterate that the formalism of Li:2023wlo goes beyond the local effective Lagrangians that exact energy–momentum conservation in the process is, by its very nature, based upon, due to the presence of Finsler dynamical background (34), where foam defects act as an external quantum-gravity ‘environment’, resulting in energy violations for particle reactions Ellis:2000sf . As we discussed in Li:2023wlo , for a $1\rightarrow 2$-body decay process (relevant for us here), the induced energy nonconservation in the observable particle subsystem, is associated with the kinetic energy of the heavy D-particle in the recoiling foam: [^14] For a $\cal N$}-particle interaction, a theorem holds, stating, that the total energy of all the observable particles is in general violated by the requirement of energy conservation for the complete matter ++ recoiling D-brane system: $\prescript{({\cal N})}{}{\sum\nolimits_{\ell}^{\pm}}E_{\ell}\simeq\frac{M_{s}}{2g_{s}}u_{\|({\cal N})}^{2},\quad\prescript{({\cal N})}{}{\sum\nolimits_{\ell}^{\pm}}{\bf p}_{\ell}=M_{D}{\bf u}_{\|({\cal N})},\quad\Rightarrow\quad\delta\overline{E}_{D}^{({\cal N})}\coloneqq\overline{\prescript{({\cal N})}{}{\sum\nolimits_{\ell}^{\pm}}E_{\ell}},$ here $u_{\|({\cal N})}$ is a proper generalization of (34), c.f. ${\bf u}_{\|}=\sum_{\ell}^{\pm}({\bf p}_{\ell}/M_{D})$ and the $\sum^{\pm}$ represents, e.g., in (42): $\sum^{\pm}\!E_{\ell}=E_{\textrm{in}}-\sum E_{\textrm{out}}=E_{1}-E_{2}-E_{3}$ as ${\cal N}=3$. For details see appendix A of Li:2023wlo and/or ref. Ellis:2000sf .
@@
$2m_{e}^{2}=E_{th}\Bigl(4\delta\overline{E}_{D}^{(3)}-\frac{g_{s}}{2M_{s}}\mathfrak{d}_{\gamma}^{2}E_{th}^{2}\Bigr),\quad E_{th}\coloneqq E_{e,th}^{\textrm{D-foam}}$ (45)
@@
where according to Li:2023wlo we assume this equality to hold. The condition, that the relativistic electron/positron of momentum $\bf p$}, responsible for producing UHE IC radiation of the Crab Nebula, is near or below the corresponding threshold implied by (45),
@@
$\frac{g_{s}}{2M_{s}}\bigl(2\varsigma_{{\cal I}}{}^{\gamma}-\mathfrak{d}_{\gamma}^{2}\bigr)\lvert{\bf p}\rvert^{3}\leq 2m_{e}^{2}.$ (48)
@@
We have to stress that these values imply no extra suppression of the effect, but natural values for both parameters in such brane models where both $\mathfrak{d}_{\gamma}^{2}$ and $\varsigma_{{\cal I}}{}^{\gamma}$ being dimensionless though, are free to vary, independently. In case a lower (TeV) $M_{D}$ is considered (as in some low-scale string schemes Antoniadis:1999rm ) the scenario envisaged by the data requires a smaller variance; this can also be naturally realized, with a much more dilute population of defects, provided $\mathfrak{d}_{\gamma}^{2}$ is proportional to the D-particle density, $n_{D}$, which is another free parameter of the bulk string cosmology model of Ellis:2004ay we used here. Although, from the observational requirement one concludes a limit (49) for $\mathfrak{d}_{\gamma}^{2}>2\varsigma_{{\cal I}}{}^{\gamma}$, this is not fine-tuning in our opinion, but implies the ‘recovery’ of special-relativistic kinematics, which is indeed respected by strings, due to underlying Lorentz symmetry that is broken, or better, reduced by the (stringy) QG ground state. Indeed we have noted that the model may be characterized Li:2023wlo via a reduced-Lorentz symmetry (akin to the one found from coupling Ellis:1999sf a fermionic system to the gravitational field (24)), with which the challenges brought about by superluminal neutrino instabilities, are properly resolved. These have not been discussed in our previous study of this isotropic stochastic D-foam, presented in Li:2023wlo , where attention was restricted to neutrino speed with favorable signal of the effect in astrophysical data.
@@
While our analysis relies exclusively on the microscopic language of conservation laws, before closing this section, we briefly comment on its relationship to macroscopic approach. Classically, whether a moving electron radiates depends on whether the signal fronts of the radiation emitted at different points along its trajectory can interfere constructively in the far field. The speed that matters for the coherence condition ${\bf v}_{e}\cdot{\bf k}=\omega({\bf k})$ (which reduces in a LI vacuum with $\omega=kv_{ph}=k$ to $v_{e}\cos\theta_{C}=v_{ph}=1$, where $\theta_{C}$ is the Cherenkov angle) is the phase velocity $v_{ph}(k)=\omega/k$, determined by the dispersion relation. The alert reader should note that in the stretched-string model previously discussed, the apparent slowdown of photon propagation arises from nonlocal effects (that is, global delays (25)) which do not alter the phase matching for local interference. Since the photon dispersion relation remains unchanged in this theory, $\overline{\omega}/\overline{k}\neq 1/\eta_{D}(\overline{\omega})$, even if $v_{e}$ exceeds the effective average speed of photons delayed by D-particle interaction, $v_{e}<1$ still leads to destructive interference of the radiated waves, aligning with observations at the quantum level. Within models featuring a random foam, however, the phase speed of light may not be retained, potentially offering a classical explanation for vacuum Cherenkov effect under specific conditions. However, the analysis of the front velocity $v_{fr}=v_{ph}(\infty)$ requires a nonperturbative analysis going beyond the approximations made above for that case; this remains therefore an open question for further investigation.
@@ 4 Evasion of Cherenkov and other strict limits on photons
We know that modified dispersions of the type (1) if used to encapsulate QG phenomenology have a broad scope to limit LV. Besides vacuum Cherenkov effect focused on here, other effects regarding photon sector include time-of-flight delay in $\gamma$-ray detections. In Shao:2009bv ; Li:2020uef in particular, it was suggested that, even with the incorporation of intrinsic timing of $\gamma$-rays at the respective source a tiny subluminal speed variation linearly scaling as photon energy could still be a significant contribution to the observed time delay, with an associated scale $E_{\textrm{LV}\gamma}\simeq 3\times 10^{17}$ GeV for the effect $\Delta v_{\gamma}=-E_{\gamma}/E_{\textrm{LV}\gamma}$. If one attempts to interpret it with the corresponding anomalous dispersion relations of the photons, the hard Cherenkov effect is allowed by kinematics as a result. It is in that context that UHE electrons deduced from the LHAASO measurements of the Crab IC spectrum have been claimed to raise the scale well above the Planck energy He:2022jdl ; specifically, the presence of 2.3 PeV long living electrons requires the scale to be greater than $1.2\times 10^{25}$ GeV, which exceeds the sensitivity of GRB photon time delay studies to such subluminal photon LV properties Shao:2009bv ; Li:2020uef by about seven orders of magnitude.
@@
[…]
@@
[…]
@@
We must emphasize once more, however, that any speculation about Lorentz violations in cosmic radiation needs support from more data. With many more UHE astrophysical $\gamma$-ray data being accrued by LHAASO and other experiments, it might be able to disentangle source from possible propagation effects due to fundamental (string) physics. On the other hand, from the perspective of the D-particle models endorsed here, other cosmic probes like neutrinos may help to discriminate between the above two approaches: the string-stretched linear-in-energy effects (30), (31), when applied to neutrinos, stay the same Ellis:2008gg , while for the models of Li:2023wlo they become entangled with the particle/antiparticle nature of neutrinos as evidenced from eq. (35). A more detailed recapitulation of their respective predictions can be found elsewhere Li:2025stf .
@@ 5 Discussion
In this paper we have revisited electron Cherenkov effect in vacuo, a nonstandard Lorentz-breaking effect for which we have no evidence till today, and thus, whose absence in particular has been used by some recent analyses He:2022jdl ; Li:2022ugz to improve prior limits on certain types of Lorentz violations in QED. Clearly, the observational hint against such a process is indeed relevant for a pure-kinematics phenomenological formalism (where just modified dispersion is postulated) as adopted therein, but in ways that it seems difficult to quantify, due to the dependence on the strength of interaction (dynamics/matrix element), which is not known in this context. So, here, we have contemplated the implications of these findings from the perspective of the comments offered in the introduction section 1 concerning the plausibility (and, practically, lack thereof) of conspiracies between modifications to kinematics and that of dynamics. We see that the latter is not particularly important as deriving the rates, and it is verified that in cases of our interest, above the threshold the rate is so tremendous that, it would allow a threshold analysis alone with trans-Planckian sensitivity which relies on photon observations from distant nebulae such as the Crab.
@@
However as it has become clear in section 4, one cannot use this result to set constraints to certain models of (stringy) QG, according to which effective subluminal refractive indices of a 4D nonstandard vacuum, as ‘seen’ by the photons, may be induced, via specific types of reduced-Lorentz invariance. In particular, in section 3, we discuss such instances arising in brane-world scenario and we see that within D-foam models hereby provided albeit linearly modified photon propagation is found to characterize the situation electrons and, in general charged excitations, remain undisturbed by the presence of space-time D-brane defects. At first sight this is just the LV parameter configuration of $(\eta=0$, $\xi<0)$, allowed by a general setting (1), pointing likely towards the possibility of hard Cherenkov effect. While in such noneffective field theories for QG foam, this is not the case. In subsection 3.2, we described briefly the reasoning based on two specific foam situations, including the cases of stretched strings and isotropic, stochastic recoil fluctuations. In both cases, although fermions could travel faster than photons with comparable energies, at least effectively no such a radiation happens. Therefore in the contexts of such models one can develop a consistent description for the intriguing result of refs. Shao:2009bv ; Li:2020uef where a linearly Planck-scale suppressed light-speed variation was suggested from analyzing travel times of GRB and AGN photons, and which may thus serve as a support to such string models Li:2021gah .
@@
To make our paper most useful for the reader who looks for the severest bound ever to this phenomenon, and hence on these two types of modifications in particle dispersions, we note that these scales can be further constrained with LHAASO’s latest report on an UHE $\gamma$-ray bubble LHAASO:2023uhj in the direction of the active star-forming region Cygnus-X, the region from which the previous $\sim 1.4$ PeV photon event was found. Gamma-rays obtained towards the Cygnus bubble are observed up to an unprecedented high energy of 2.48 PeV. [^17] We note that the detected UHE photons spread over the entire bubble region make unlikely the leptonic IC origin of the radiation and indicate the involvement of hadronic process LHAASO:2023uhj , so that the argument leading to electron Cherenkov bound as applied to the Crab is not valid in this case. So, we find the constraints:
@@
The estimates we make tighten the strongest one (among the existing values, say of refs. LHAASO:2021opi ; Li:2021tcw ; He:2022jdl ) by a factor of $\sim(12/7)^{3}\approx 5$. The inclusion of more general cases having $\xi\neq 0$ and $\eta\neq 0$, is straightforward, and, parallels the analyses carried out in He:2022jdl using 1.4 PeV as the greatest photon energy, prior to the discovery of the $\sim 2.5$ PeV event. While an analysis in this line does not present any technical difficulty, it requires a separate investigation which takes us out of the purpose of the present paper.
Finally, one should observe that the trans-Planckian lower limits for these types of LV, or equivalently, the small values of the corresponding dimensionless parameters $\xi$ and $\eta$ are consistent with the subluminal feature of the D-foam-induced vacuum refractive indices for photons, capable of accommodating the hint of light-speed variation exposed in the time-of-flight tests Shao:2009bv ; Li:2020uef . Indeed in the context of such brane models, as advocated here, photons are stable (i.e., do not decay), so that Li:2021gah the strong $\gamma$-decay constraints thereof LHAASO:2021opi ; Li:2021tcw ; He:2022jdl , as well as the most stringent result just derived, are inapplicable. For reconciliation of such models with LHAASO data, especially in this respect, we refer the reader to our paper Li:2021tcw for details. Moreover, it should be noted that the peculiarity of the transparency of D-foam to electrons or charged probes, for specifically stringy properties outlined above, [^18] It is for the same reason that such models also survive the severe constraints (see, e.g., He:2024ljr ) that would be obtained from models with LV modifications in protons and/or pions and thus modifications in hadronic GZK cutoff, which is unaffected due to vanishing or suppressed interactions between particles involved (i.e., protons, or pions whose constituents are charged) and D-foam. provides additional reasons for escaping the otherwise very stringent bounds for putative LV effects on electrons/positrons, including (55). For this latter point, see also Li:2022ugz . This is similar to the way, discussed in that paper, out of electron (soft) Cherenkov limits on this parameter imposed from the positive side.